OpenStudy (anonymous):

Giving medel! The first term of an arithmetic sequence is -5, and the tenth term is 13. Find the common difference.

OpenStudy (blurbendy):

how many units do you have to go to get from -5 to 13?

OpenStudy (anonymous):

+18

OpenStudy (blurbendy):

oh, wait, first term is -5, tenth is 13

OpenStudy (blurbendy):

so if you went, -5, -3, -1, 1, 3, 5, 7, 9, 11, 13, this will give you 13 as your tenth term

OpenStudy (blurbendy):

so what's the common difference?

OpenStudy (anonymous):

2

OpenStudy (blurbendy):

yup, good job

OpenStudy (whpalmer4):

You want to divide the difference between the final term and the starting term by the difference of the term numbers. \[\frac{13-(-5)}{10-1} = \frac{18}{9} = 2\]

OpenStudy (blurbendy):

cool @whpalmer4